Opened 7 years ago

Closed 4 years ago

Last modified 15 months ago

#457 closed flight processing (fixed)

RG12/10, flight day 250/2012, Luton Bedford

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors: edfi besm

Description (last modified by besm)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/RG12_10-2012_250_Luton_Bedford

Data arrived from ARSF via SATA disk on 24/09/2012

RG12/10 Scientific objective:

RG12/10 Priority: a4h

PI: Karen Anderson

Notes: Lidar (incl. FW) is of the higest priority, followed by RCD, followed by HS)

Sensors:

Camera (30/01/2013)
Eagle (30/01/2013)
Hawk (30/01/2013)
LiDAR (30/01/2013)
FW Lidar (30/01/2013)

Change History (38)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)

comment:2 Changed 6 years ago by knpa

  • Description modified (diff)
  • Summary changed from RG12/10, flight day 250a/2012, Luton Bedford to RG12/10, flight day 250/2012, Luton Bedford

comment:3 Changed 6 years ago by mark1

Base station data has been resent by Ops - original appeared to be corrupt - this one may be better

comment:4 Changed 6 years ago by besm

Navigation has been processed. I have completed two versions:

  • Navigation data as usual (good quality solution)
  • Navigation data based on basestation data at a 30 second interval (good quality solution)

This is to attempt processing with each and see whether there is a big difference between data processing with the full basestation data and the 30-second basestation data. Relevant folders in the posatt/ subdirectory are basestation_30s/ and ipas_honeywell_30s/.

comment:5 Changed 6 years ago by besm

Started LiDAR processing and found these pitch and roll errors:

Pitch: -0.003054378596439879
Roll: 0.0003787571920568071

comment:6 Changed 6 years ago by besm

Completed RCD processing and readme. Ready to delivery check.

comment:7 Changed 6 years ago by besm

Finished classification of discrete LiDAR points. Could now go on to create delivery with discrete LiDAR data, or wait for full-waveform processing to become viable. Suspending LiDAR processing until decision is made.

comment:8 Changed 6 years ago by edfi

Starting RCD check

comment:9 Changed 6 years ago by edfi

RCD check complete all ok

comment:10 Changed 6 years ago by besm

Started processing LiDAR FWD.

comment:11 Changed 6 years ago by besm

While checking FWD, found pitch+roll values to mismatch wildly. It turns out the first and second group of flightlines seem to have radically differing pitch+roll values. Correcting now.

comment:12 Changed 6 years ago by knpa

Renamed Eagle and Hawk raws so they match and are sequential (and got rid of .1 bits caused by crash).

Eagle 8 is missing (the file numbered 8 is just 30 seconds captured in-between flightlines).

Hawk 19 and 20 are missing (from Hawk dying for this period - see logsheet)

comment:13 Changed 6 years ago by knpa

SCTs

e h
1 -0.05 -0.06
2 -0.05 -0.03
3 -0.03 -0.06
4 -0.05 -0.06
5 -0.03 -0.06
6 -0.02 -0.06
7 -0.06 -0.06
8 n/a -0.06
9 -0.06 -0.06
10 -0.05 -0.06
11 -0.01 -0.04
12 -0.05 -0.06
13 -0.02 -0.04
14 -0.03 -0.05
15 -0.05 -0.06
16 -0.05 -0.06
17 1.34 -0.06
18 -0.02 -0.06
19 -0.02 n/a
20 -0.03 n/a
21 -0.03 -0.05
22 -0.01 -0.03
23 -0.01 -0.02
24 -0.06 -0.06
25 -0.02 -0.06
26 -0.05 -0.06
27 -0.02 -0.06

Eagle 17 did not have any nav information

comment:14 Changed 6 years ago by knpa

  • Other processors set to edfi besm
  • Owner set to knpa
  • Status changed from new to assigned

comment:15 Changed 6 years ago by besm

Found pitch and roll to vary a few times. Found estimations as follows.

Start Flightline End Flightline Pitch Roll
131942 134645 -0.003074 0.0003146
135211 145159 -0.002980 0.0003805
145611 145611 -0.002980 0.0004933
150522 154020 -0.003054 0.0003788

comment:16 Changed 6 years ago by edfi

Hyperspectral delivery check completed

comment:17 Changed 6 years ago by besm

Completed LiDAR which should be ready to delivery check now.

comment:18 Changed 6 years ago by anhi

LiDAR Delivery Check:

  • Moved FW document into delivery.

No other problems found.

comment:19 Changed 6 years ago by besm

Delivered all sensors to Karen Anderson+Antoine Cottin on 30/01/2013 on disks 148+150.

comment:20 Changed 6 years ago by besm

  • Description modified (diff)

comment:21 Changed 6 years ago by edfi

Starting vector check for LiDAR and hyperspectral

comment:22 Changed 6 years ago by edfi

Vector check complete ready for re-check

comment:23 Changed 6 years ago by anhi

LiDAR vector delivery check:

  • Needs to be a FW LiDAR readme.
  • Needs to have a vector accuracy estimate so the readme knows the screenshots are vector overlaid.

Screenshots look good

Last edited 6 years ago by anhi (previous) (diff)

comment:24 Changed 6 years ago by anhi

HS vector delivery check:

No problems found.

comment:25 Changed 6 years ago by besm

Ammended the issues discussed:

  • replaced .trj files
  • remade readme with vectors

comment:26 Changed 6 years ago by anhi

FW LiDAR Deliver check:

"An initial QC of your data shows that it agrees with the Ordance Survey vectors on average to
within 0-2 meters." repeated in the readme under Data quality and remarks heading.

No other problems.

comment:27 Changed 6 years ago by besm

Should be fixed now.

comment:28 Changed 6 years ago by anhi

FW LiDAR Delivery check:

No Problems

comment:29 Changed 5 years ago by emca

  • Component changed from Processing: general to Archiving

comment:30 Changed 5 years ago by emca

Archiving this project

Removed *bil + *bil.hdr files (lines 01-05) from delivery directory - they were there in addition to the zipped files.

comment:31 Changed 5 years ago by emca

As mentioned above base station data was processed twice.
Not sure which base station data was used to process the lidar so leaving them both. There are two reg files pointing to different base station solutions.

comment:32 Changed 5 years ago by emca

Dem has incorrect name (aster.dem). Leaving as is, to avoid confusion later.

comment:33 Changed 5 years ago by emca

Ready for archiving

Email sent to Wendy 15/01/2014

comment:34 Changed 4 years ago by dac

Reprocessed LiDAR lines 23 and 27 in ALSPP for Steve Hancock, exporting AGC in Flight ID field and sent via FTP on 12/02/2015

comment:35 Changed 4 years ago by dac

Reprocessed all LiDAR lines for Steve Hancock, exporting AGC. Sent via FTP on 06/02/2014

comment:36 Changed 4 years ago by dac

Archiving

Started uploading data to NEODC

comment:37 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Archived data available at NEODC - looks like old version without AGC value saved in full waveform LAS files. Will notify Wendy but marking as closed.

Last edited 4 years ago by dac (previous) (diff)

comment:38 Changed 15 months ago by asm

Archiving

Reprocessed LiDAR data has been uploaded. JSON files created and uploaded as well and CEDA has been notified (31/01/2018).

Note: See TracTickets for help on using tickets.